Commissioning Notes 20080417 UT

Observers: J. M. Hill, R. M. Wagner
Telescope Operator: A. Ceranski

Summary

There was very little observing due to high winds. We were open for 80 minutes in bad seeing, but lost half of that to an OSS problem. The sky was photometric, but winds were gusting above 25 m/sec all the rest of the night. We did get some software tested, and we got the position angle set through a zeropoint in PCS.

Dome Closed Activities

(Times in UT)

Other Telescope Activities

We were able to donate a couple hours to James so he could get Laser Tracker measurements of the swing arm positions as a function of telescope elevation.

Aaron had problem with DX hardpoint 0 chasing to the TI limit during a raise. In hindsight we thing it was because the TCS on lbtmu101 had lost contact with the other servers.

Several cell-related computers in the lower treehouses had to be rebooted after UPS power in that area was off for a while today.

Collimation Model Updates

I've created SXSMIRTCFCollimation.dat for the Fixed Secondary. It has a global offset for the nominal (almost centered) hexapod position. But there is no position dependence with elevation or temperature. This file is in the new location of /home/telescope/TCS/Configuration/PSF/.

I've also analyzed the combined IRTC collimation data from the pointing stars from 20080413 and 20080415 combined. This has been installed as SXPMIRTCFCollimation.dat. I made an improved version of mkcoll.pro which compensates the collimation values for temperature. There was a 3.5 degC temperature change during 20080415. By adjusting the Z-values for the known temperature coefficient of +0.1 mm/degC I was able to reduce the residuals from ~200 microns rms to ~80 microns rms. Spurred on by this success and 50 mph winds outside, I also found by eye temperature coefficients for the X and Y position. The result is less clear, but the residuals were reduced from ~300 microns rms to ~200 microns rms. The X-coefficient was -0.08 mm/degC and the Y-coefficient was -0.14 mm/degC. This collimation model is about equal to the best we have been able to do with LBC. However, armed with this knew tool of adjusting the temperature coefficient for each point in the fit, we can probably reprocess the old LBC data and make a better model. The main uncertainty has to do with temperature gradients. Temperature gradients in the telescope during the first half of 20080415 could have simulated equilibrium temperature effects in X and Y. Only more data over a range of temperatures will tell for sure. There are now separate versions of mkcoll for LBC and for IRTC since the Elevation value is radians in the IRTC headers and degrees in the the LBC headers.

The data, the collimation file, and some GIF plots live in the usual place of /home/lbcobs/Collimation/

Intra/Extra Cube Pre-Processing

Mark has finished the IRAF script to pre-process IRTC data cubes for feeding extra/intra pairs into EF. I guess all we need now is some data.

PCS Rotator Tests

06:20 We tested the new PCS software by editting the PCSInstrument.conf file in its new location in
/home/telescope/TCS/Configuration/PCS/PCSInstrument.conf

I've editted the file to make LEFTZEROPOINT 45.0 (degrees).
This should rotate the polynomials sent to MCSPU so that IRTC is properly rotated to PA=0, but I'm not sure about the sign of the offset. (+135 turned out to be a better answer.)

I've also made irtc.20080415.ptmod the default pointing model for IRTC at BGF.

Mark thought of the clever launcher command to have ds9 start with /Repository as the default directory: csh -c "cd /Repository; ds9"

Some Actual Observing

08:00 Wind dying down to merely 10 m/sec.

08:08 Open shutter doors

08:17 pointto WT10_262, but didn't find a star.

irc is reporting "Invalid buffer RPC error". We figured out that this was coming from OSS. Stopped and restarted OSS several times, even on a different machine. Went to the treehouse to check on UMAC (that we know nothing about). Found a dangling network cable and plugged it in, but this turned out to be the M3 UMAC. Where is the M2 UMAC? Decided we had to call Paul and wake him up. Paul says the RPC error is from OSS talking to MCSPU (about swing arm status). We stop and restart MCSPU and that gets rid of the RPC error, but OSS is still complaining about Invalid Buffer. M2 UMAC responds to "ver" query, but appears not to respond to init command. Finally Paul kicks OSS in the right way (to be debugged further when he is here tomorrow.) and hexapod is alive again.

08:58 back in business on the sky, but no star

9:01 spiral search, found star at offset 20,-15 with the old collimation model.

Initialize to new collimation model and search again. Found star at -10,30 and collimation is worse. Needed to add SX M1 global offset of X=0.1 Y=0.6 Z=0.9. This Y correction is cancelling the Y temperature term that I built into the collimation model earlier. Of course, the telescope is ~9 degC out of equilibrium, so miscollimation is not so surprising. Seeing is ~4 arcsec.

09:15 Change Rotator offset to -45. This makes N down and E left at PA=0.

09:20 Preset to PA=180. This makes N up and E right.

Back to PA=0 and Change Rotator offset to +135 deg. N is now up as expected for PA=0. And N is down for PA=180, so rotator seems working appropriately.

09:30 Close dome for increasing wind. Open two vent doors halfway to cool down the chamber.

09:45 Testing the irc OffsetGuiding command. It will guide with "irc OffsetGuiding 0.0 1.0 0.0 RADEC_FOCAL left", but PCS makes the 1 arcsec move in azimuth rather than RA.

10:20 Close the vent doors. Mirror Ventilation continues to run to cool down the mirrors.

Jose has changed irs to give filenames in UT, so now well ask Stephen to change the prefix configuration to "irtc." which will give us an LBT standard filename.

10:45 Wind continues

-- JohnHill - 17 Apr 2008
Topic revision: r4 - 20 Apr 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback